# Toward expert-level medical text validation with language models

### v4

@@ 6 Conclusion
In this research, we introduce MedVAL, a novel, self-supervised distillation method for training LMs to validate LM-generated medical text. Our benchmark highlights substantial improvements in model performance and provides systematic evidence that LMs can achieve performance statistically non-inferior to a single human expert ($p<0.001$). The MedVAL-Bench dataset, with its diverse tasks and physician annotations, establishes a robust benchmark for validating LM-generated medical text. To support further research and adoption, we release: 1) Codebase, 2) MedVAL-Bench, and 3) MedVAL-4B (the best-performing open-source model). Together, these contributions lay a foundation for developing and deploying safer, more trustworthy tools to automate the process of validating LM-generated medical text, reducing physician burden. The findings from this study suggest that incorporating MedVAL into clinical workflows could enhance the reliability of medical text generation, thereby improving patient safety and clinical efficiency. Future research should explore prospective studies to validate the practical benefits of MedVAL in real-world settings, potentially leading to an integrated approach that can support healthcare professionals and enhance patient care.
@@ 7 Data and Code Availability
    Codebase for fine-tuning and evaluation: github.com/StanfordMIMI/MedVAL
@@
    MedVAL-4B (best-performing fine-tuned Qwen3-4B): huggingface.co/stanfordmimi/MedVAL-4B
Datasets used in this study include MedicationQA [47], MeQSum [48], Open-i [49], MIMIC-IV [50], MIMIC-IV-BHC [51] and ACI-Bench [52, 53, 54]. All these datasets (except MIMIC) are fully open-source. The MIMIC-IV datasets require PhysioNet access and are currently undergoing review to be made available on PhysioNet. Further distribution of datasets is subject to the data sharing agreements stipulated by the original creators.
@@ 8 Acknowledgements
We acknowledge the support of the Advanced Research Projects Agency for Health (ARPA-H) Chatbot Accuracy and Reliability Evaluation (CARE) project, as well as helpful discussions around stakeholder needs and technical solutions guided by program managers, Mansoor Khan, and Dr. Berkman Sahiner.
A.A. is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. A.C. receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, A.C. receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, and Brain Key; E.A. receives consulting fees from Fourier Health.
This work was supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the Advanced Research Projects Agency for Health (ARPA-H).
@@ 9 Author Contributions
AA collected data, developed code, ran experiments, designed studies, analyzed results, created figures, and wrote the manuscript. All authors reviewed the manuscript, providing meaningful revisions and feedback. VB, MV, NC, SO, AS, MP, AK provided specific feedback on the manuscript and technical advice. AJ, KAM, EJPG, PNR, SG, CB, EPR, EDZVR, PLH, KRK, MG, EL, DBL participated in the reader study, with AJ, EPR, SG, CB, and EDZVR serving as radiologists, while the others participated as internal medicine physicians. EJPG, KAM, and PNR participated as bilingual internal medicine physicians. DBL, CL, RD, JH, SK advised on study design and provided feedback. EA and ASC guided the project and advised on technical details. RD and ASC obtained funding for this project and managed the overall project. No funders or third parties were involved in study design, analysis, or writing.
Appendix A Physician Reader Study Instructions
@@ conclusions.
Appendix B Generator Prompt
@@
Appendix C Validator Prompt

### v5

@@ 6 Conclusion
We introduce MedVAL, a novel, self-supervised distillation method for training LMs to validate LM-generated medical text. Our benchmark highlights substantial performance improvements and provides evidence that LMs can achieve performance statistically non-inferior to a single human expert on a subset annotated by multiple physicians ($p<0.001$). The MedVAL-Bench dataset, with its diverse tasks and physician annotations, establishes a benchmark for validating LM-generated medical text. To support further research and adoption, we release: 1) Codebase, 2) MedVAL-Bench, and 3) MedVAL-4B (the best-performing open-source model). Together, these contributions lay a foundation for developing and deploying safer, more trustworthy tools to automate the process of validating LM-generated medical text, reducing physician burden. The findings from this study suggest that incorporating MedVAL into clinical workflows could enhance the reliability of medical text generation, thereby improving patient safety and clinical efficiency. Future research should explore prospective studies to validate the practical benefits of MedVAL in real-world settings, potentially leading to an integrated approach that can support healthcare professionals and enhance patient care.
@@ 7 Data and Code Availability
    MedVAL-Bench (PhysioNet) [46]: physionet.org/content/MedVAL-Bench
  4. 4.
  5. 5.
    MedVAL Demo: stanfordmimi/MedVAL
Datasets used in this study include MedicationQA [48], MeQSum [49], Open-i [50], MIMIC-IV [51], MIMIC-IV-BHC [52] and ACI-Bench [53, 54, 55]. All these datasets (except MIMIC) are fully open-source. Further distribution of datasets is subject to the data sharing agreements stipulated by the original creators.
8 Funding
This research was, in part, funded by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Government. Furthermore, this work was supported in part by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the ARPA-H.
AA is supported by NIH grant R01 HL167974 and ARPA-H contract AY2AX000045. ASC receives research support from NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060; ARPA-H contracts AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. Unrelated to this work, ASC receives research support from GE Healthcare, Philips, Microsoft, Amazon, Google, NVIDIA, Stability; has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, and Brain Key; EA receives consulting fees from Fourier Health.
@@ 9 Acknowledgements
We acknowledge the support of the ARPA-H Chatbot Accuracy and Reliability Evaluation (CARE) project, as well as helpful discussions around stakeholder needs and technical solutions guided by program managers, Mansoor Khan, and Dr. Berkman Sahiner.
@@ 10 Author Contributions
AA collected data, developed code, ran experiments, designed studies, analyzed results, created figures, and wrote the manuscript. All authors reviewed the manuscript, providing meaningful revisions and feedback. VB, MV, NC, SO, AS, MP, and AK provided specific feedback on the manuscript and technical advice. AJ, KAM, EJPG, PNCR, SG, CB, EPR, EDZVR, PLH, KRK, MG, EL, and DBL participated in the reader study, with AJ, EPR, SG, CB, and EDZVR serving as radiologists, while the others participated as internal medicine physicians. EJPG, KAM, and PNCR participated as bilingual internal medicine physicians. DBL, CL, RD, JH, and SK advised on study design and provided feedback. EA and ASC guided the project and advised on technical details. RD and ASC obtained funding for this project and managed the overall project. No funders or third parties were involved in study design, analysis, or writing.
11 Competing Interest
No competing interests to declare.
Appendix S1 Additional Discussion and Limitations
Our formulation combines absolute and relative generator-validator consistency as a soft filtering signal rather than a hard constraint. While relative consistency is algebraically implied by absolute calibration under ideal conditions, in practice, validator scores are noisy and not perfectly calibrated. The absolute terms capture calibration on clean and perturbed outputs, while the relative term encourages sensitivity to perturbation-induced changes between paired generations. Intuitively, this difference term acts as a regularizer on the absolute scores, favoring examples where the validator reflects meaningful factual degradation even when absolute scores may share bias or drift. While our implementation uses discrete degradation, future work can explore the role of relative consistency when validator outputs represent continuous factual degradation.
MedVAL is designed to distill an explicit, scalar notion of factual degradation conditioned on the input, rather than to learn a general-purpose embedding space or rely on surface-level similarity. Contrastive representation-learning approaches and overlap-based metrics such as ROUGE or BERTScore primarily capture lexical or embedding-level similarity between outputs, which may remain high even when clinically meaningful errors are introduced (e.g., incorrect recommendation). As a result, such metrics are not well aligned with risk-calibrated medical text validation, where factual consistency must be assessed relative to the input and clinical semantics. Exploring representation-based or contrastive objectives as alternative validator training strategies remains an interesting direction for future work.
We adopt zero-shot CoT prompting as a reasonable and widely used baseline for inference-time adaptation. We focus on contrasting this setting with lightweight self-supervised fine-tuning, as it provides a clear and controlled comparison between prompting-only adaptation and weight updates. We leave a systematic analysis of automatic prompt optimization methods on MedVAL tasks as an important avenue for future work.
While discharge summaries and BHC narratives are primarily clinician-facing artifacts, clinically dense English documentation from the hospital course is often adapted into patient-facing discharge instructions and after-visit summaries. These patient-facing materials are routinely provided in multiple languages in U.S. hospital systems, and prior work has shown that linguistically concordant discharge materials improve patient activation and clinical outcomes for individuals with limited English proficiency [82]. Because Spanish is among the most commonly requested languages, we include English-to-Spanish transformation of hospital course-derived narratives as a high-information, cross-lingual stress test. Our goal is not to evaluate translation systems or simulate exact documentation workflows, but to assess whether MedVAL can reliably detect factual inconsistencies, omissions, or clinically meaningful certainty shifts when dense content is transformed.
The datasets used in MedVAL-Bench originate from academic medical centers, which may not fully capture the heterogeneity of real-world clinical environments. To probe robustness under distribution shift, we evaluate MedVAL-distilled models on MEDEC, a separate benchmark for medical error detection, and observe consistent improvements over zero-shot baselines. Furthermore, we emphasize that MedVAL is intended as a proof-of-concept benchmark and validation framework, rather than a deployment-ready system. Future work should expand evaluation to data from real-world clinical workflows to further assess external validity.
Our study does not evaluate MedVAL within a prospective human-AI workflow or assess downstream clinical impacts; MedVAL is intended as a decision-support tool for risk triage rather than a replacement for physician oversight. A prospective validation study could be conducted by deploying MedVAL alongside LM-generated clinical documentation in real-world workflows and measuring its agreement with physician assessments on consecutively generated cases, including its ability to prioritize high-risk outputs for review.
Importantly, the goal of MedVAL-distilled open-source models is not to surpass proprietary frontier models, but to enable small and open-source validators to achieve competitive performance that supports low-cost, privacy-preserving, and scalable deployment in clinical settings. Further, our non-inferiority analysis is restricted to the multi-annotated subset (15 examples $\times$ 6 tasks; 90/840). It should be interpreted as evidence of expert-aligned performance under the evaluated risk-based validation setting, rather than as a general claim of expert equivalence across all tasks or clinical contexts.
Appendix S2 Physician Reader Study Instructions
@@ conclusions.
Appendix S3 Generator Prompt
@@
Appendix S4 Validator Prompt